OpenStudy (anonymous):

Help me pleaseee . fan and medal Solve the following system of equations: 2x + 3y − z = 1 3x + y + 2z = 12 x + 2y − 3z = −5 (3, 1, 2) (−3, 1, 2) (3, −1, 2) (3, 1, −2)

OpenStudy (anonymous):

@hero @myininaya @dan815

OpenStudy (anonymous):

@aaronq

OpenStudy (anonymous):

@hartnn @kropot72

OpenStudy (redneckoutlaw):

C I believe

OpenStudy (anonymous):

How'd you get C?

OpenStudy (redneckoutlaw):

7x + 7y = 14 -(11x + 7y = 26) ~ -4x = -12 - - -> x=3 11(3) + 7y = 26 33 + 7y = 26 - - -> y=-1 2x + 3y - z = 1 2(3) + 3(-1) -z = 1 6 + (-3) -z = 1 3-z = 1 -z = 1-3 -z = -2 - - -> z=2